Find the domain of the following:
1.f(x)=4.5e
2. g(x)=2^x Show your work and explain how you obtained your answer
3. h(x)=-e^x
Description of transformation
Equation(s) for the Horizontal Asymptote(s).
y-intercept in (x,y) form:

Answer by stanbon(48535)

Find the domain of the following:
1.f(x)=4.5e
Domain: all Real Numbers
Description: horizontal line thru all points where y = 4.5*2
Horizontal Asymptote: none
y-intercept:
Let x = 0, then y = 4.5e
intercept is (0,4.5e)
------------------------------
2. g(x)=2^x Show your work and explain how you obtained your answer
Domain: all Real Numbers
Description: increasing exponential function
Horizontal Asymptote: y = 0
y-intercept: Let x=0, then y = 1
Intercept is (0,1)
------------------------------
3. h(x)=-e^x
Domain: all Real Numbers
Description: decreasing exponential function
Horizontal Asymptote: y = 0
y-intercept: Let x = 0, then y = -1
Intercept is (0,-1)
===================================
Cheers,
Stan H.